OpenStudy (anonymous):

Solve for x. x2 + 9x = 0 A. 0, -9 B. 0, 9 C. 1, -9 D. 1, 9

OpenStudy (rmrjr22):

x(x+9) = 0

OpenStudy (rmrjr22):

x= 0 then also x+9 = 0

OpenStudy (rmrjr22):

then just solve

OpenStudy (rmrjr22):

get it?

OpenStudy (anonymous):

so the answer is b ?

OpenStudy (rmrjr22):

no solve for x, get x alone

OpenStudy (anonymous):

i dont have to solve the problem i have to pick an answer !

OpenStudy (rmrjr22):

but thats how you get an answer, i promise its easy

OpenStudy (rmrjr22):

solve for x... to do that you must get it alone. To do that you need to separate x from the equation to eliminate x^2 and from the 9.

OpenStudy (anonymous):

its okay i moved on from that question

OpenStudy (rmrjr22):

so it becomes x(x+9) = 0 .. then from there set both values of = to 0, so that way ie. x+9 = 0.... you can move 9 to the other side (subtract it) and what do you get? X all by its self... which gives you 0 and -9 as the values of x.....

OpenStudy (rmrjr22):

Best response please?